# EcoLoRA: Communication-Efficient Federated Fine-Tuning of Large Language Models

### v1

@@ 5 Conclusion
In this paper, we introduced EcoLoRA, a novel communication-efficient federated fine-tuning framework for LLMs. Our approach comprises a round-robin segment sharing scheme, an adaptive sparsification method, and lossless encoding. Extensive evaluations on QA and VA tasks across diverse datasets and models show that EcoLoRA substantially reduces communication overhead while maintaining accuracy. Moreover, it remains robust under non-i.i.d. settings and incurs minimal computational overhead.
@@
Due to constraints in computational resources and time, we focused on a single, representative federated learning setting, following standard practice in existing studies. In future work, we plan to explore multiple federated learning scenarios involving diverse client participations to further validate the generalizability of our approach. Additionally, we restricted our experiments to Llama models for question-answering tasks and the Vicuna model for value alignment tasks; investigating a wider range of models will be crucial for fully demonstrating the robustness and versatility of our method.
@@ Appendix A Experimental Settings
To simulate non-i.i.d. data distribution across clients, we divide the datasets using a Dirichlet distribution with $\alpha=0.5$. For the Dolly dataset, we directly use the provided category labels for splitting. Since the Alpaca dataset lacks explicit categories, we generate synthetic ones. Specifically, we concatenate the ‘instruction’ and ‘input’ fields of each sample into a single string, convert these strings into TF-IDF vectors (using up to 1000 features and excluding English stop words), and apply KMeans clustering to group samples based on textual similarity. The resulting clusters are treated as synthetic categories, and client-specific datasets are created by applying a Dirichlet-based allocation to these clusters. Additionally, we consider a more heterogeneous non-i.i.d. scenario in which each client is assigned data from a distinct task domain.
@@ Appendix B Convergence Proof
$=-\frac{\eta}{2}(1-\eta L)\|\nabla F(P_{t})\|^{2}-\frac{\eta}{2}(1-\eta L)\|E_{t}\|^{2}$
@@
$\mu\|\nabla F(P_{t})\|^{2}\leq F(P_{t})-F(P_{t+1})+\eta\left(2\eta L-1\right)\cdot\Delta$
Summing both sides over $t=0$ to $T-1$:
$\sum_{t=0}^{T-1}\mu\|\nabla F(P_{t})\|^{2}\leq F(P_{0})-F^{*}+T\eta\left(2\eta L-1\right)\cdot\Delta$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(P_{t})\|^{2}\leq\frac{F(P_{0})-F^{*}}{\mu T}+\frac{\eta\left(2\eta L-1\right)\Delta}{\mu}$
@@ Appendix C Additional Ablation Study
Threshold $k$ Fixed Top-$k$ Adaptive Sparsification
0.9 66.5 66.6
0.7 66.1 66.5
0.6 66.1 66.5
0.5 65.8 66.3

### v2

@@ 5 Conclusion
In this paper, we introduced EcoLoRA, a novel communication-efficient federated fine-tuning framework for LLMs. Our approach comprises a round-robin segment sharing scheme, an adaptive sparsification method, and lossless encoding. Extensive evaluations on QA and VA tasks across diverse datasets and models show that EcoLoRA substantially reduces communication overhead while maintaining accuracy. Moreover, it remains robust under non-IID settings and incurs minimal computational overhead.
Acknowledgment
We thank the reviewers for their valuable feedback. This work was partially supported by NSF (CNS-2154930, CNS-2229427, CNS-2238635, CCF-2403758), ARO (W911NF-24-1-0155, W911NF-25-1-0059), and ONR (N00014-24-1-2663, N00014-24-1-2730).
@@ Acknowledgment
The primary limitation of this work is that EcoLoRA is developed and evaluated only with LoRA. Although LoRA is currently the most widely adopted PEFT method for federated fine-tuning of LLMs, this focus limits the broad applicability of our method.
On the other hand, we believe the core design principles of EcoLoRA are broadly applicable to other PEFT methods as well. For instance, adapter-based approaches introduce modular trainable layers between frozen backbone parameters, and prefix-tuning leverages independent prefix vectors prepended to transformer blocks. Both of these methods exhibit modularity and layer-wise independence, making them naturally compatible with EcoLoRA’s segment-sharing strategy. We leave a systematic exploration of applying EcoLoRA to these and other PEFT methods to future work.
@@ Appendix A Experimental Settings
To simulate non-IID data distribution across clients, we divide the datasets using a Dirichlet distribution with $\alpha=0.5$. For the Dolly dataset, we directly use the provided category labels for splitting. Since the Alpaca dataset lacks explicit categories, we generate synthetic ones. Specifically, we concatenate the ‘instruction’ and ‘input’ fields of each sample into a single string, convert these strings into TF-IDF vectors (using up to 1000 features and excluding English stop words), and apply K-means clustering to group samples based on textual similarity. The resulting clusters are treated as synthetic categories, and client-specific datasets are created by applying a Dirichlet-based allocation to these clusters. Additionally, we consider a more heterogeneous non-IID scenario in which each client is assigned data from a distinct task domain.
@@ Appendix B Convergence Proof
$=-\frac{\eta}{2}(1-\eta L)\|\nabla F(P_{t})\|^{2}$
$\quad-\frac{\eta}{2}(1-\eta L)\|E_{t}\|^{2}$
@@
$\mu\|\nabla F(P_{t})\|^{2}$ $\leq F(P_{t})-F(P_{t+1})$
$\quad+\eta\left(2\eta L-1\right)\cdot\Delta$
Summing both sides over $t=0$ to T1T-1:
$\sum_{t=0}^{T-1}\mu\|\nabla F(P_{t})\|^{2}$ $\leq F(P_{0})-F^{*}$
$\quad+T\eta\left(2\eta L-1\right)\cdot\Delta$
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(P_{t})\|^{2}$ $\leq\frac{F(P_{0})-F^{*}}{\mu T}$
$\quad+\frac{\eta\left(2\eta L-1\right)\Delta}{\mu}$
@@ Appendix C Additional Ablation Study
Client Selection Strategy. In realistic FL scenarios, clients have heterogeneous hardware and network conditions, making uniform sampling relatively less representative. To this end, we evaluate EcoLoRA under asynchronous FL, where clients update at different speeds. Following FedBuff Nguyen et al. (2022), the server aggregates updates from a buffer of the fastest clients, allowing them to participate more frequently. Results with LLaMA2-7B on Dolly (Table 10) confirm that EcoLoRA maintains strong performance even under asynchronous client participation.
Table 10: Comparison of accuracy and parameters (in millions) under different client selection strategies.
Method ARC Upload P. Total P.
FedIT 66.3 3523.2 6962.5
w/ EcoLoRA 66.4 508.6 3955.8